# Benchmarking digital quantum simulations and optimization above hundreds of qubits using quantum critical dynamics

### v1

@@ Acknowledgments
We thank Julien Gacon, Elena Peña Tapia, and Almudena Carrera Vazquez for help with the code, and Stefan Wörner, Laurin Fischer, Samuele Piccinelli, and Francesco Tacchino for insightful and inspiring discussions. G. M. acknowledges financial support from the Swiss National Science Foundation (grant PCEFP2_203455). This research was supported by the NCCR MARVEL, a National Centre of Competence in Research, funded by the Swiss National Science Foundation (grant number 205602). We acknowledge the use of IBM Quantum services for this work. The views expressed are those of the authors, and do not reflect the official policy or position of IBM or the IBM Quantum team.
Appendix A Error mitigation and suppression
@@
A.1 Readout error mitigation
@@ A.2 Dynamical decoupling
A.3 Pulse-efficient transpilation
@@
Appendix B Hardware properties and qubit selection
[…]
./figures/res_en_statevector_one_column.pdf
@@ Appendix B Hardware properties and qubit selection
Appendix C Quantum annealing with very large time steps
[…]
./figures/kink_kink_correlator.pdf
@@ Appendix C Quantum annealing with very large time steps
Appendix D Kink-kink correlator
Since the KZ scaling of the density of defects could be explained classically [86], we investigate here the correlation between different defects in spin chains after digitized quantum annealing, also termed kink-kink correlation in the literature [30]. Given uniform couplings $J$, the solution after annealing is ferromagnetic. In this setting, defects are misalignments of spins on edges $i$ between lattice sites, and the correlator between defects $i$ and $i+r$ is defined as
@@ Appendix D Kink-kink correlator
Fig. 8 shows the density of defects (left) of annealing a periodic 12-qubit chain on ibm_auckland (a) using REM and pulse-efficient transpilation and of a 100-qubit chain on ibm_sherbrooke (b) using no EMS. The corresponding kink-kink correlators for different $t_{\mathrm{f}}$ are shown in the respective right panel as a function of the normalized lattice distance $r/\xi$ with $1/\xi=n_{\mathrm{def}}$. The positive peaks between $r/\xi=0.5$ and $1.0$ imply quantum fluctuations responsible for the defect production.

### v2

@@ Acknowledgments
We thank Julien Gacon, Elena Peña Tapia, and Almudena Carrera Vazquez for help with the code, and Stefan Wörner, Laurin Fischer, Samuele Piccinelli, and Francesco Tacchino for insightful and inspiring discussions. We also thank Adolfo Del Campo, Jacek Dziarmaga, and Andrew King for providing useful feedback. G. M. acknowledges financial support from the Swiss National Science Foundation (grant PCEFP2_203455). This research was supported by the NCCR MARVEL, a National Centre of Competence in Research, funded by the Swiss National Science Foundation (grant number 205602). We acknowledge the use of IBM Quantum services for this work. The views expressed are those of the authors, and do not reflect the official policy or position of IBM or the IBM Quantum team.
Appendix A Influence of hardware noise
Figure 6: Density of defects scaling in the presence of simulated hardware noise. a All curves show simulated Trotterized time evolution of a 12-qubit chain with $\Delta t=0.5$. The noise-free reference curve (pink line) represents a statevector simulation. The hardware curve (purple diamonds) corresponds to results obtained from ibm_sherbrooke using only REM. All other curves show results of sampling $10^{5}$ measurements from a simulated circuit and include simulated hardware noise that corresponds to the full noise model of ibm_sherbrooke. b Error of the density of defects relative to the noise-free results.
The effects of noise such as a finite temperature or bath-couplings on QA have been previously studied both numerically [63, 28] and experimentally [31] in analog settings. The result is a deviation from the KZ scaling and a subsequent increase in defect density.
Here, we perform simulations of the Trotter algorithm including a tunable noise model. Moreover, since noise generally leads to decoherence of the system, this further validates our claim that observing a density of defects that follows the KZ scaling implies coherent evolution. Scaling the average device errors across several orders of magnitude allows for a detailed understanding of their influence on the density of defects. In other words, the noisier the system, the sooner we expect decoherence and, in turn, a deviation from the expected scaling.
The noise model is constructed from error rates of ibm_sherbrooke obtained through standard calibrations [72]. To simplify the scaling of the noise, we use error rates obtained from the average of individual gate and qubit properties. Concretely, our model applies the following noise to all qubits,
$T_{1}=T_{1}^{\mathrm{ave}}/\eta\ ,\quad T_{2}=T_{2}^{\mathrm{ave}}/\eta$ (13)
$e_{\mathrm{1q}}=\eta e_{\mathrm{1q}}^{\mathrm{ave}}\ ,\quad e_{\mathrm{2q}}=\eta e_{\mathrm{2q}}^{\mathrm{ave}}\ ,\quad e_{\mathrm{ro}}=\eta e_{\mathrm{ro}}^{\mathrm{ave}}\ .$
[…]
Fig. 6a shows the density of defects obtained from QA of a periodic 12-qubit system with simulated noise scaled by $\eta\in\{0.01,\ldots,10\}$. The noise-free line is a statevector simulation of the same system. In addition, we plot real hardware results of simulating the same 12-qubit system on ibm_sherbrooke, using only REM. The results of this experiment match the noisy simulations with unscaled noise $\eta=1$, thereby validating our choice of noise model. Both the noisy and the noise-free simulations as well as the hardware experiment were done with the same time step of $\Delta t=0.5$ as in Section III.
Since the system size that we can simulate exactly with noise is small, the behavior of the density of defects over time appears less smooth compared to larger sizes. These finite size effects manifest as a drop in the noise-free density of defects after roughly $t_{\mathrm{f}}\approx 10$. The noise-free simulations indicates a monotonically decreasing density of defects with $t_{\mathrm{f}}$. In the noisy cases instead, after a threshold time $t_{\mathrm{f}}^{*}$, the density of defects starts increasing with time. As the noise strength decreases, $t_{\mathrm{f}}^{*}$ increases. This is further underlined by Fig. 6b, showing the relative error of the density of defects with respect to the noise-free curve, $|(n_{\mathrm{def}}(\eta)-\tilde{n}_{\mathrm{def}})/\tilde{n}_{\mathrm{def}}|$, with $\tilde{n}_{\mathrm{def}}$ the noise-free defect density. While this behavior is compatible with earlier numerical studies on dissipative quantum annealing and analog hardware experiments [63, 31, 28], the main difference here is that the density of defects always keeps increasing with time after $t_{\mathrm{f}}^{*}$. In contrast, dissipative quantum dynamics or annealing experiments [31, 32], indicated that the density of defects may decrease again, possibly with the same $-1/2$ exponent, even beyond the coherence limit.
Appendix B Trotter error of benchmarking experiments
Figure 7: Density of defects scaling for different discretizations $\Delta t$. Statevector simulations of a periodic 12-qubit chain comparing the exact, i.e., non-discrete time evolution to Trotterized simulations with different time steps $\Delta t$. The time step $\Delta t=0.5$ is highlighted in pink and corresponds to the time step chosen for the hardware experiments in Section III.
Discretization of the time evolution requires setting a time step, which in turn controls the discretization error, in our case the Trotter error. Ideally, the time step should be sufficiently small to reduce the Trotter error and accurately describe the continuous-time limit of the dynamics. However, smaller time steps require more Trotter circuit layers to reach the same final time. In Section III, we adopt a practical approach and utilize a small, yet non-vanishing, time step of $\Delta t=0.5$.
Fig. 7a shows statevector simulations of a periodic 12-qubit chain with time steps varying between $\Delta t=0.01,\ldots 1.0$ and compares the resulting density of defects to the exact, continuous-time result. The figure shows that our choice of $\Delta=0.5$, despite minor deviations from the continuum solution, provides a stable evolution over the entire time range we consider. This is not the case for larger time steps as can be seen by the increased fluctuations. Crucially, this time step allows us to experimentally probe relevant regimes of $t_{\mathrm{f}}$. More concretely, the KZ scaling we seek to benchmark against is not present for small annealing times $t_{\mathrm{f}}$, and observing the KZM requires probing sufficiently large $t_{\mathrm{f}}$. However, this is not a limitation of the proposed method. Rather, it is a limitation of the current hardware but as the hardware improves and longer circuits, i.e., more Trotter steps, can be reliably executed, smaller time steps can be adopted to improve the accuracy at smaller $t_{\mathrm{f}}$.
Appendix C Error mitigation and suppression
@@
C.1 Readout error mitigation
@@ C.2 Dynamical decoupling
C.3 Pulse-efficient transpilation
@@
Appendix D Hardware properties and qubit selection
@@
Appendix E Quantum annealing with very large time steps
@@
Appendix F Kink-kink correlator
We investigate the correlation between different defects in spin chains after digitized quantum annealing, also termed kink-kink correlation in the literature [31, 109]. Given uniform couplings $J$, the solution after annealing is ferromagnetic. In this setting, defects are misalignments of spins on edges $i$ between lattice sites, and the correlator between defects $i$ and $i+r$ is defined as
@@ Appendix F Kink-kink correlator
Fig. 10 shows the density of defects (left) of annealing a periodic 12-qubit chain on ibm_auckland (a) using REM and pulse-efficient transpilation and of a 100-qubit chain on ibm_sherbrooke (b) using no EMS. The corresponding kink-kink correlators for different $t_{\mathrm{f}}$ are shown in the respective right panel as a function of the normalized lattice distance $r/\xi$ with $1/\xi=n_{\mathrm{def}}$. The existence of a positive peak, which we observe between $r/\xi=0.5$ and $1.0$, is expected from theoretical results[81, 68, 110].